# Advancing Compositional Awareness in CLIP with Efficient Fine-Tuning

### v1

@@ 4.4 Generalization to larger architectures and models
Next, we measure the generalization of the method to larger architectures such as ViT-L/14 across models pre-trained on different datasets (CLIPA and CLIPS). Results on ViT-B/16 can be found in Table 16 in the appendix. We fine-tune each of the pre-trained models, CLIP, CLIPA, and CLIPS, on our 1M Laion subset and on RedCaps (PixelProse) subset using CLIC. As most other compositionality enhancing methods do not explore larger architectures, the only baselines we have are CoN-CLIP and our NegCLIP† (the details of which can be found in Appendix D). In Table 3, we can see that for all types of pre-trained models, CLIC-RedCaps shows consistent improvements on the ITT score for SugarCrepe++: +7.4% for CLIP, +9.4% for CLIPA, and +9.2% for CLIPS when averaged across replace and swap. These improvements are on a similar scale as the ones for the ViT-B/32 model, suggesting that even for larger architectures, CLIC can attain substantial gains.
@@
What we find even more fascinating are the improvements for image and text retrieval attained by CLIC. Specifically, for CLIPS (which has to the best of our knowledge SOTA retrieval numbers of a CLIP model), CLIC further improves text retrieval by 1.3% and image retrieval by 2.2% on MS-COCO. These gains comes at a tiny fraction (0.01%) of the pre-training cost, as we only see 850k (RedCaps) resp. 1M (Laion) samples compared to 13B samples seen during pre-training of CLIPS.
4.5 Contrasting against SOTA
In Table 4, we pitch our best model in CLIPS + CLIC against the SOTA CLIP like and other models on SugarCrepe++. To the best of our knowledge, CLIPS + CLIC attains the best available ITT numbers on SugarCrepe++ across all architectures. We note that CLIPS in itself is already good for ITT on SugarCrepe++ but our short fine-tuning enhances it significantly further in terms of compositional understanding. For TOT, among CLIP like models CLIPS + CLIC performs similarly to the best model (CLIPS) in this category. We were unable to evaluate BLIP-Large for TOT, since the individual BLIP-Large encoders are not publicly available. The WinoGround results for CLIPS + CLIC are the best in text and group score, while the image score is the best among CLIP-like models. As far as we know, CLIPS + CLIC attains SOTA WinoGround scores among CLIP based models.
@@ 5 Conclusion
We introduce CLIC, a simple, low-cost, and scalable method for boosting compositionality in vision-language models, as measured by SugarCrepe++. This improvement comes with a minimal degradation in zero-shot classification, but enhanced text and image retrieval performance. CLIC works across different good caption quality datasets. We further illustrate the generalization of CLIC to variedly (dataset/training-scheme) pre-trained models and architectures. Lastly, we show that CLIC even improves performance on models already exhibiting strong compositional and retrieval abilities like CLIPS and yields SOTA results on SugarCrepe++ as well as image retrieval.
@@ Acknowledgements
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting AP and NDS. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy (EXC number 2064/1, project number 390727645). We are also thankful for the support of Open Philanthropy and the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors.
Appendix A …Limitations
@@
Appendix B …Dataset preparation
@@
Appendix C …Algorithmic details
@@
Appendix D …Experimental details and discussions
@@
Appendix E …Additional experiments
@@ Appendix A Limitations
Our work considers only a fraction of contrastive models, we believe CLIC should work similarly for other models like SigLIP [50], which we were unable to test due to compute and time constraints. Since all gains attained by CLIC are due to fine-tuning the text encoder, we cannot test CLIC fine-tuned models in large LVLMs like LlaVA [28] or VILA [26], which only use the vision encoder from CLIP.
@@ Appendix B Data preparation
In this section, we describe the data preparation process in detail. Appendix B.1 showcases randomly selected images and captions from the PixelProse [44] dataset and from the Laion [41] dataset, along with captions generated for these images by CogVLM [47]. This part also presents further details on these datasets. In Section B.2, we explain the hard-negative generation procedure. In Section B.3 we show examples of the limitations of using a single image for generating negatives as done in [49]. Section B.4 shows ablations regarding other ways of combining images for CLIC.
@@ B.1 Training datasets
In this section, we present our data generation process. To show that CLIC performs well across diverse datasets, we train models from three different sources. We took a subset of images from Laion with captions that were generated by CogVLM [47], using the query presented in Figure 3. We also used a recently proposed high-quality caption dataset in PixelProse [44], which contains available samples from the RedCaps subset. In order to compare faithfully to our baselines in Table 1 who use CC3M, we also use a 1.2M subset of CC12M that comes from the same data stream as CC3M to train a ViT-B/32 model. Since our method requires detailed captions, we use the recaptioned version of CC12M from PixelProse. These two datasets come from a different stream of internet data than Laion. A selection of random image-caption pairs from both kinds of datasets is provided in Figure 4. As can be seen from the examples, the first sentence in each caption usually provides a general description of the image, while the other part of the caption often captures more specific details. To clean the data, we removed the starting strings like “This picture depicts/shows/demonstrates:”. Then, since the generated captions tend to be longer than those in the standard datasets like Laion and DataComp (the average caption length in our recaptioned dataset is 78 words), we split the captions into individual sentences as done in [30, 51]. This results in an average sentence length of 16 words for captions generated by CogVLM [47]. We retain only images whose captions contain more than one sentence, which is more than of the 1M set for Laion. This approach favors captions that provide both general context and specific details about the images. For a fair comparison with NegCLIP, we also train CLIC with MS-COCO, the details of which are in Section 4.
@@ B.3 In-scene vs across-scene swaps
Swapping words within a single sentence can lead to negatives due to the changes in word order. For example, the sentence from NegCLIP: “The horse is eating the grass" can become “The grass is eating the horse". However, swapping words across sentences from different images greatly increases the diversity of negative examples. Taking the example from the paper (Section 1), concatenating the image with an image of a dog chasing a cat may lead to the new negative: "The horse is chasing the cat. The dog is eating the grass". Since each image can be paired with each of the other images, this scales the diversity of negatives quadratically and helps the generalization of the method.
@@
Creating negatives from a single scene, like in NegCLIP, may result in a modified sentence that still accurately describes the image. This might happen as in many sentences, changing the order of the words does not change the semantic meaning of the sentence. This is different from changing words across scenes, where the concatenated sentence will reflect the concatenated image only if we by chance swap words that have the same meaning (note that we do not allow swaps of the same word, i.e., ‘man’ can not be swapped with ‘man’). Here are a few examples from the CC12M subset from PixelProse:
@@ B.4 Non-random image concatenation
Combining images at random offers us several advantages such as computational efficiency and reducing the overfit to specific benchmarks. However, to test whether using related scenes might produce harder negatives and, by that, help us achieve better results (at the expense of higher computational resources) we combine images that share a common noun. To this end, we first extract all the nouns from the first sentence of each image using spaCy. Then for each image, we sample up to five images (without repetition) that share a common noun and use this to concatenate the images. We then continue as usual by swapping random words (excluding the common noun). Although we witnessed a higher loss (Eq. 5), indicating that the hard negatives are indeed harder, the results on the downstream tasks stayed roughly the same, as can be seen in Table 6
@@ Appendix C Algorithmic details
In this section, we present a pseudo-code of our algorithm, as discussed in Section 3 in the main paper. The pseudo-code can be found in Algorithm 1.
@@ D.1 Further training details and discussions
As stated in the main paper, we train with alternating steps of CLIP loss (Eq. 1) and the proposed loss (Eq. 7). This decision is made to help the fine-tuned model retain its performance on downstream tasks, evident from the ablations in Table 10.
Computational resources.
All the work was carried out on A100 40G GPUs. The training runs per hour are across 4 GPUs. Running CLIC on smaller architectures like ViT-B/32 and ViT-B/16 took less than 30 minutes per GPU hour for the Laion dataset. Datasets such as RedCaps took longer due to high-resolution images. The larger ViT-L/14 models took around 1 GPU hour per dataset. The evaluation time is negligible, in the order of 20 minutes per model for all datasets on a single GPU.
Training parameters and details.
In the experiments, the loss parameters in Eq. 7 are set to , and . We train for one epoch on our 1M Laion subset and on our PixelProse dataset, while for MS-COCO, we trained for five epochs; ablation on the number of epochs can be found in Table 11. All experiments are conducted at an image resolution of . Our effective batch size is , and we use a cosine scheduler, where the warm-up phase is of the training time. The learning rate (LR) starts at , peaks at , and arrives at at the end. We used the standard AdamW [31] optimizer with beta parameters and set to with a weight decay of 0.1.
Training with COCO.
Although for most of the experiments, we use our 1M Laion subset or the PixelProse dataset, we also train with MS-COCO to be comparable to the original NegCLIP checkpoint from [49]. Note that due to training on MS-COCO all retrieval numbers on MS-COCO for NegCLIP from [49], and CLIC-COCO are no longer zero-shot. The same is also true to CoN-CLIP, who use COCO images as distractor in their method.
Our NegCLIP baselines.
To have a baseline for all models considered in this work, we train our own version of NegCLIP (denoted everywhere with †). It is used only when there are no available checkpoints (Table 3, Table 16 and Table 17), and in our ablation study (Table 12). This version of NegCLIP differs from the NegCLIP [49] as follows:
@@
We use their negative loss calculation (see Eq. 2), where the hard negatives are part of the standard clip loss.
@@ D.3 Dataset from other methods
To improve compositionality, many methods use re-labelled/captioned versions of standard datasets. The details of these recaptions can be found in Table 7. TripletCLIP uses recpationed data from LaCLIP and DAC uses BLIP2 [23] to create better captions. Methods like CoN-CLIP and SVLC-R+L use the original captions and PaLM-2 and BERT [6] to generate hard-negatives for their base datasets.

### v2

@@ 4.4 Generalization to larger architectures and models
Next, we measure the generalization of the method to larger architectures such as ViT-L/14 across models pre-trained on different datasets (CLIPA and CLIPS). Results on ViT-B/16 can be found in Table˜24 in the appendix. We fine-tune each of the pre-trained models, CLIP, CLIPA, and CLIPS, on our 1M Laion subset and on RedCaps (PixelProse) subset using CLIC. As most other compositionality enhancing methods do not explore larger architectures, the only baselines we have are CoN-CLIP and our NegCLIP† (the details of which can be found in Appendix˜D). In Table˜3, we can see that for all types of pre-trained models, CLIC-RedCaps shows consistent improvements on the ITT score for SugarCrepe++: +7.4% for CLIP, +9.4% for CLIPA, and +9.2% for CLIPS when averaged across replace and swap. These improvements are on a similar scale as the ones for the ViT-B/32 model, suggesting that even for larger architectures, CLIC can attain substantial gains.
@@
What we find even more remarkable are the improvements for image and text retrieval attained by CLIC. Specifically, for CLIPS (which has to the best of our knowledge SOTA retrieval numbers of a CLIP model), CLIC further improves text retrieval by 1.3% and image retrieval by 2.2% on MS-COCO. These gains come at a tiny fraction (0.01%) of the pre-training cost, as we only see 850k (RedCaps) resp. 1M (Laion) samples compared to 13B samples seen during pre-training of CLIPS.
Constrasting again SOTA. In Table˜5, we pitch our best model in CLIPS + CLIC against the SOTA CLIP like and other models on SugarCrepe++. To the best of our knowledge, CLIPS + CLIC attains the best available ITT numbers on SugarCrepe++ across all architectures. We note that CLIPS in itself is already good for ITT on SugarCrepe++ but our short fine-tuning enhances it significantly further in terms of compositional understanding. For TOT, among CLIP like models CLIPS + CLIC performs similarly to the best model (CLIPS) in this category. We were unable to evaluate BLIP-Large for TOT, since the individual BLIP-Large encoders are not publicly available. The WinoGround results for CLIPS + CLIC are the best in text and group score, while the image score is the best among CLIP-like models. As far as we know, CLIPS + CLIC attains SOTA WinoGround scores among CLIP based models.
4.5 CLIC vision encoder in LLaVA
Unlike our previous experiments, we fine-tune the CLIP model with CLIC at resolution while unfreezing both the vision and text encoders. We then replace the CLIP vision encoder in LLaVA-1.5-7b (Vicuna) [31] with our CLIC version and fine-tune the full LLaVA model following standard projector pre-training (full details in Section˜D.5). As shown in Table˜6, CLIC enabled LLaVA-1.5 achieves notable compositional reasoning improvements. For SugarCrepe++, we observe gains across most categories with marginal degradation for swap object, while WinoGround scores improve for text and group sets by more than 6% and 3% respectively. Crucially, these gains preserve performance on standard VLM benchmarks (GQA, TextVQA, SQA-I), with MME (perception) improving from 1441.0 to 1465.3. These results show that even short compositional fine-tuning via CLIC can significantly enhance fine-grained visual-linguistic understanding while maintaining general vision-language capabilities, as further illustrated by qualitative examples in Figure˜6.
| SugarCrepe++ (ITT) | WinoGround | VLM-tasks | ||||||||||
| Model | Replace | Swap | ||||||||||
| Obj | Att | Rel | Obj | Att | Text | Img | Grp | GQA | TextVQA | SQA-I | MME (P) | |
| LLaVA-1.5 (CLIP) | 94.3 | 80.4 | 73.3 | 66.9 | 69.8 | 36.7 | 38.2 | 23.2 | 60.6 | 57.1 | 69.2 | 1441.0 |
| LLaVA-1.5 (CLIC) | 94.3 | 80.8 | 73.8 | 66.5 | 71.2 | 43.5 | 42.0 | 26.2 | 62.1 | 57.7 | 68.2 | 1465.3 |
@@ 5 Conclusion
We introduce CLIC, a simple, low-cost, and scalable method for boosting compositionality in vision-language models, as measured by SugarCrepe++. Notably, this improvement comes with a minimal degradation in downstream tasks like zero-shot classification and enhanced text and image retrieval performance. CLIC works across different good caption quality datasets. We further illustrate the generalization of CLIC to variedly (dataset/training-scheme) pre-trained models, architectures. Lastly, we show that CLIC even improves performance on models already exhibiting strong compositional and retrieval abilities like CLIPS and yields SOTA results on SugarCrepe++ as well as image retrieval. From our initial experiments, CLIC powered LLaVA improves in compositional reasoning over the standard version. We believe compositionally aware vision encoders can further help large VLMs like LLaVA on more tasks, a detailed study of this is left for future work.
@@ Acknowledgements
The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting AP and NDS. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy (EXC number 2064/1, project number 390727645), as well as in the priority program SPP 2298, project number 464101476. We are also thankful for the support of Open Philanthropy and the Center for AI Safety Compute Cluster. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the sponsors. We thank the reviewers for the discussions and suggesting additional experiments.
Appendix˜A …Limitations
@@
Appendix˜B …Dataset preparation
@@
Appendix˜C …Algorithmic details
@@
Appendix˜D …Experimental details and discussions
@@
Appendix˜E …Additional experiments
@@ Appendix A Limitations
Our work considers only a fraction of contrastive models, we believe CLIC should work similarly for other models like SigLIP [57], which we were unable to test due to compute and time constraints. Moreover, we were only able to test LLaVA [31], whereas other large VLMs like VILA [28] and CogVLM [53] might also benefit from CLIC, which we leave for future work.
@@ Appendix B Data preparation
In this section, we describe the data preparation process in detail. Appendix B.1 showcases randomly selected images and captions from the PixelProse [50] dataset and from the Laion [46] dataset, along with captions generated for these images by CogVLM [53]. This part also presents further details on these datasets. In Section˜B.2, we explain the hard-negative generation procedure. In Section˜B.3 we show examples of the limitations of using a single image for generating negatives as done in [56]. Section˜B.4 shows ablations regarding other ways of combining images for CLIC.
@@ B.1 Training datasets
In this section, we present our data generation process. To show that CLIC performs well across diverse datasets, we train models from three different sources. We took a subset of images from Laion with captions that were generated by CogVLM [53], using the query presented in Figure˜3. We also used a recently proposed high-quality caption dataset in PixelProse [50], which contains available samples from the RedCaps subset. In order to compare faithfully to our baselines in Table˜1 who use CC3M, we also use a 1.2M subset of CC12M that comes from the same data stream as CC3M to train a ViT-B/32 model. Since our method requires detailed captions, we use the recaptioned version of CC12M from PixelProse. These two datasets come from a different stream of internet data than Laion. A selection of random image-caption pairs from both kinds of datasets is provided in Figure˜4. As can be seen from the examples, the first sentence in each caption usually provides a general description of the image, while the other part of the caption often captures more specific details. To clean the data, we removed the starting strings like “This picture depicts/shows/demonstrates:”. Then, since the generated captions tend to be longer than those in the standard datasets like Laion and DataComp (the average caption length in our recaptioned dataset is 78 words), we split the captions into individual sentences as done in [33, 59]. This results in an average sentence length of 16 words for captions generated by CogVLM [53]. We retain only images whose captions contain more than one sentence, which is more than of the 1M set for Laion. This approach favors captions that provide both general context and specific details about the images. For a fair comparison with NegCLIP, we also train CLIC with MS-COCO, the details of which are in Section˜4.
@@ B.3 In-scene vs across-scene swaps
Swapping words within a single sentence can lead to negatives due to the changes in word order. For example, the sentence from NegCLIP: “The horse is eating the grass" can become “The grass is eating the horse". However, swapping words across sentences from different images greatly increases the diversity of negative examples. Taking the example from the paper (Section˜1), concatenating the “The horse is eating the grass" image with an image of a dog chasing a cat may lead to the new negative: "The horse is chasing the cat. The dog is eating the grass". Since each image can be paired with each of the other images, this scales the diversity of negatives quadratically and helps the generalization of the method.
@@
Creating negatives from a single scene, like in NegCLIP, may result in a modified sentence that still accurately describes the image. This might happen as in many sentences, changing the order of the words does not change the semantic meaning of the sentence. This is different from changing words across scenes, where the concatenated sentence will reflect the concatenated image only if we by chance swap words that have the same meaning (we remind the reader that we do not allow swaps of the same word, i.e., ‘man’ can not be swapped with ‘man’). Here are a few examples from the CC12M subset from PixelProse:
@@ B.4 Non-random image concatenation
Combining images at random offers several advantages, such as computational efficiency and reduced risk of overfitting to specific benchmarks. However, to examine whether generating more plausible or grammatically correct negatives leads to better results (at the expense of higher computational resources), we construct negatives in a non-random manner using two strategies:
1.
Combining images that share a common noun.
2.
Swapping words with different semantic meanings and selecting images accordingly.
Common nouns
We begin by extracting all the nouns from the first sentence of each image using spaCy. For each image, we then sample up to five other images (without repetition) that share a common noun and use this to concatenate the images. We then continue as usual by swapping random words (excluding the common noun). Although we witnessed a higher loss (Eq.˜5), indicating that the hard negatives are indeed harder, the results on the downstream tasks remained largely unchanged (see Table˜8, row Same Noun).
LLM swaps
To test the impact of non-meaningful or ungrammatical negatives, we made the following modification to our method to construct negatives that are grammatically correct by leveraging LLM based swaps:
1.
We extracted the fine-grained Part-of-Speech (POS) tags for each word in the first sentence using spaCy’s most computationally expensive model, which took around 5 hours. In contrast, the version used for the random image-caption pairs employed a more computationally efficient model and took only a few minutes.
2.
For each word, we queried Gemini 2.5-flash [5] to generate up to five replacement candidates from the same POS category, such that replacing the word changes the meaning of the caption, while keeping the sentence coherent. Note: these queries to Gemini via API calls took over 24 hours of compute time to find word replacements for the 1M Laion subset. In comparison, fine-tuning a CLIP model with CLIC only takes around 40 minutes.
3.
During training, for each caption, we randomly selected a POS category and then randomly chose a word from that category. The second image and word replacement were selected based on the candidates from step 2.
This significantly more expensive variant of CLIC (LLM Swaps in Table˜8) results in harder negatives (as evidenced by a higher loss during training) and leads to improvements in Replace (ITT) for SugarCrepe++ but degrades Swap (ITT) and TOT. While at first surprising, we hypothesize that the lower diversity of negatives compared to our original computationally cheap version of CLIC, with random replacement of words from the same POS category, could be a reason for this.
To check the quality of the negatives, we hand-labelled 75 negative samples from both LLM Swaps and the random negatives used to train CLIC-LAION. This process yielded 69% meaningful, grammatically correct negatives with distinct semantic meanings from LLM Swaps, compared to 28% from the CLIC-LAION. We include below 5 coherent and grammatically correct hard negatives examples, and 5 nonsensical or grammatically incorrect ones. While there are some errors, the overall quality is sufficient to prevent the model from relying solely on shortcuts, as many examples remain meaningful and grammatically correct, and the errors are relatively subtle.
Correct hard negative examples:
(postcard telegram)
N: a collage of various telegram marketing materials. a vintage postcard.
(stripes swirls)
N: a black satin jacket with white swirls on the collar and cuffs. a graphic design element that features a large red heart at the center surrounded by intricate golden stripes and patterns
(short thick)
N: a portrait of a woman with thick textured hair that has a mix of brown and blonde colors. an indoor setting possibly a café or a bakery with a man pouring a short brown liquid which appears to be chocolate onto a white countertop.
(plain fancy)
N: a woman standing against a fancy background. a scrapbooking paper pack titled ‘sunny memories’ by plain pants designs.
(corners hearts)
N: a pair of red envelopes with intricate gold embroidery on the hearts. two wooden objects that are shaped like corners.
Nonsensical or grammatically incorrect samples:
(whistle shout)
N: a red kettle with a shout attached to its spout. a man in a dramatic pose seemingly in the middle of a punch or a whistle
(made destroyed)
N: a hand-destroyed card with a floral design. a scene of destruction with a building that appears to have been damaged or made.
(printed removed)
N: a black bag with white lines and the word ‘thule’ removed on it. one of the nuts is shown with its cap printed revealing the coiled thread inside.
(setting breaking)
N: this image captures a moment in an outdoor park breaking where two individuals are engaged in a playful activity. a bear’s face that appears to be setting through a shattered glass surface.
(appears disappears)
N: it disappears to be a motorized treadmill with a digital display on the top. the headline reads ‘german billionaire appears on matterhorn’ and it is dated april 11 2018 at 11:35 am cdt.
@@ Appendix C Algorithmic details
In this section, we present a pseudo-code of our algorithm, as discussed in Section 3 in the main paper. The pseudo-code can be found in Algorithm˜1.
@@ D.1 Further training details and discussions
As stated in the main paper, we train with alternating steps of CLIP loss (Eq.˜1) and the proposed loss (Eq.˜7). This decision is made to help the fine-tuned model retain its performance on downstream tasks, evident from the ablations in Table˜4.
Computational resources. All the work was carried out on A100 40G GPUs. The training runs are across 4 GPUs. Running CLIC on smaller architectures like ViT-B/32 and ViT-B/16 took less than 30 minutes for the Laion dataset. Datasets such as RedCaps took longer due to high-resolution images. The larger ViT-L/14 models took around 1 hour per dataset. The evaluation time is negligible, in the order of 20 minutes per model for all datasets on a single GPU.
Training parameters and details. In all of the experiments, the loss parameters in Eq.˜7 are set to , and . We train for one epoch on our 1M Laion subset and on our PixelProse dataset, while for MS-COCO, we trained for five epochs; ablation on the number of epochs can be found in Table˜19. All experiments are conducted at an image resolution of . Our effective batch size is , and we use a cosine scheduler, where the warm-up phase is of the training time. The learning rate (LR) starts at , peaks at , and arrives at at the end. We used the standard AdamW [34] optimizer with beta parameters and set to with a weight decay of 0.1.
Training with COCO. Although for most of the experiments, we use our 1M Laion subset or the PixelProse dataset, we also train with MS-COCO to be comparable to the original NegCLIP checkpoint from [56]. Note that due to training on MS-COCO all retrieval numbers on MS-COCO for NegCLIP from [56], and CLIC-COCO are no longer zero-shot. The same is also true to CoN-CLIP, who use COCO images as distractor in their method.
Our NegCLIP baselines. To have a baseline for all models considered in this work, we train our own version of NegCLIP (denoted everywhere with †). It is used only when there are no available checkpoints (Table˜3, Table˜24 and Table˜25), and in our ablation study (Table˜20). This version of NegCLIP differs from the NegCLIP [56] as follows:
@@
We use their negative loss calculation (see Eq.˜2), where the hard negatives are part of the standard clip loss.
@@ D.3 Dataset from other methods
To improve compositionality, many methods use re-labelled/captioned versions of standard datasets. The details of these recaptions can be found in Table˜9. TripletCLIP uses recpationed data from LaCLIP and DAC uses BLIP2 [25] to create better captions. Methods like CoN-CLIP and SVLC-R+L use the original captions and PaLM-2 and BERT [7] to generate hard-negatives for their base datasets.